A satellite in low Earth orbit is 225km above the surface. What is its orbital velocity. Mass of earth= 6 x 10^24 kg. Radius of earth = 6.38 x 10^6m
Answer #1

radius of orbit , r = radius of earth + height

r = 225 + 6380 = 6605 km

speed of satellite , v = sqrt(G * m/r)

v = sqrt(6.673 *10^-11 * 6 * 10^24/(6605 *10^3))

v = 7785.7 m/s

the speed of the satellite is 7785.7 m/s